OpenStudy (anonymous):

let f(x)=7x+2 and g(x)=2x^2-x-1 whats f(3)+g(3)??

OpenStudy (anonymous):

\[f(3)+g(3)=7(3)+2+2(3)^2-3-1=21+2+18-3-1=23+14=37\]